Using the perimeter given, find the dimensions of the rectangle with the greatest area using the whole units.
1. P=10
2. P=40
3. P-64
(etc.)
(This is 6th grade math homework).
Thank you
Gail (student's mom)

Answer by Alan3354(69443)

Using the perimeter given, find the dimensions of the rectangle with the greatest area using the whole units.
1. P=10
2. P=40
3. P-64
-------------
A square has the maximum area of rectangles with a specified perimeter. Do you need a proof of that?
-----------
1. P=10
s = 10/4 = 2.5
Area = s^2 = 6.25 sq units
-------------
2. P=40
s = 10
Area = 100
--------------
3. P=64
s = 16
Area = 256